### v1

# Criticality at work: scaling in the mouse cortex enhances performance

Abstract
The critical brain hypothesis posits that neural systems operate near a phase transition, optimising the transmission, storage, and processing of information. While scale invariance and non-Gaussian dynamics — hallmarks of criticality — have been observed in brain activity, a direct link between criticality and behavioural performance remains unexplored. Here, we use a phenomenological renormalisation group (PRG) approach to examine neuronal activity in the primary visual cortex of mice performing a visual recognition task. We show that non-trivial scaling in neuronal activity is associated with enhanced task performance, with pronounced scaling observed during successful task completion. When rewards were removed or non-natural stimuli presented, these scaling signatures diminished. These results suggest that critical dynamics in the brain may be crucial for optimising behavioural outcomes, offering new insights into the functional role of criticality in cortical processing.
keywords:
brain criticality, renormalisation group, task performance, cortical dynamics
Introduction
The idea that living systems self-organise in ways that promote efficiency, resilience, and adaptation has captivated scientists across multiple fields [1, 2, 3, 4, 5]. In neuroscience, this notion is encapsulated by the critical brain hypothesis [6, 7, 8, 9, 10, 11, 12, 13], which posits that the brain operates near a phase transition, a tipping point between two different states. This regime would be able to maximise information transmission, information storage and dynamic range in response to incoming stimuli [14, 15, 16, 7, 17, 18, 19, 20, 21], leading to debates about the possible clinical implications of the hypothesis [22, 23, 24, 25, 26].
One of the hallmark features of the critical point where a phase transition occurs is scale invariance: the system’s activity has no characteristic length or time. Over two decades ago, this property was initially found in the self-similar temporal structure of human electroencephalogram (EEG) data [27] and in the scale-free nature of neuronal avalanches — bursts of electrical activity — in cortical slices [6]. Since then, a growing body of evidence supporting the existence of the putative phase transition has been gathered in different experimental setups [28, 29, 30, 31, 32, 33, 34, 35].
Recently, a novel phenomenological renormalization group (PRG) method was introduced [36, 37], offering a fresh approach to studying scale-invariant behaviour in complex systems. Inspired by the renormalization group (RG) — the gold standard for understanding phase transitions in theoretical models [38, 39, 40, 41] — the PRG similarly reveals non-trivial scale invariance at non-Gaussian fixed points. It does so by tracking how observables evolve across scales as the system undergoes recursive coarse-graining. However, unlike the RG, the PRG is model-free and directly applicable to high-dimensional data from complex systems, making it particularly useful in neuroscience, where experimental advancements now enable the simultaneous recording of thousands of neurons [42]. The effectiveness of the PRG method has been validated through various studies, ranging from neural activity in C. Elegans to the whole human brain [37, 43, 44, 45, 46].
Despite the increasing methodological sophistication in the methods to assess scale invariance in neuronal data, a glaring issue remains. Namely, that a link between criticality and task performance — the alleged reason why critical dynamics should be found in the brain in the first place — has never been found. Here we bridge this gap.
Results
Since our goal is to understand the relation between brain criticality and behaviour, we analysed neuron spiking activity from the primary visual cortex (V1) of mice while they performed a visual task [47] (Fig. 1a-c). This activity was subdivided into 30 s windows (Methods 1). For each window, we calculated the covariance matrix $C_{ij}$ comprising all pairs of recorded neurons. By diagonalising this matrix we obtained a set of eigenvectors from which projectors were created. When projecting the activity of each neuron over all eigenvectors, one recovers the original time series. By progressively selecting fewer components corresponding to the largest eigenvalues for projection, we generated an increasingly coarse-grained representation of the neural activity (a procedure akin to discarding faster modes in momentum space in standard RG) [36] (Fig. 1d). We tracked the normalised distribution of this activity as we chose fewer covariance eigenvectors. This procedure creates a flow of activity distributions (Fig. 1e-f) that pushes non-critical behaviour towards a trivial (Gaussian) fixed point. For each window, we used the kurtosis $\kappa$ of the distribution in the final coarse-graining step as a metric to quantify the deviation from the Gaussian distribution.
Figure 1: The phenomenological renormalisation group applied to mouse brain data. a, In the first part (Active) of the Allen Brain Observatory Visual Behavior Neuropixels experiment [47], natural images are shown in 250-ms flashes and the mouse needs to recognise a change of images in order to receive a reward. In sequence, the screen shows non-natural images such as Gabors, along with black and white flashes and a grey background. In the third part (Passive), the natural images are repeated in the same order as in the first part, but the mouse can no longer lick to be rewarded. b, Frontal slice of a mouse brain with the primary visual cortex (V1) highlighted in bright green (left) (credits to Allen Institute for Brain Science https://atlas.brain-map.org) and scheme of a Neuropixels 1.0 probe (right) used in the data collection. c, Raster plot of spiking activity for a 30-s window. d, Progressive coarse-graining of a unit time series, corresponding to projections onto a decreasing number of eigenvectors of the covariance matrix ($N,N/2,N/4,...$). The variance of the activity is normalized to one. e, Schematic representation of the renormalisation flow created by the coarse-graining procedure. Time windows that are further from criticality, such as those from spike-shuffled surrogate data, converge more rapidly to the Gaussian fixed point. f, In the final step of the coarse-graining procedure, windows closer to criticality (blue) retain their non-Gaussian shape, with higher kurtosis, while others (red) come closer to a Gaussian distribution (dashed line), with lower kurtosis.
Certain time windows exhibit clear signatures of non-trivial dynamics, as expected at a critical point, with kurtosis values significantly higher than those in the spike-shuffled (surrogate) data. In contrast, other windows show results similar to surrogate data (Fig. 1f). Across 62 sessions (from 44 mice), we observed a broad range of kurtosis values, overlapping with the distribution from the surrogate data (Fig. 2a). As we depart from the Gaussian distribution, the distinction between windows with trivial and non-trivial statistics is arbitrary. We exploit this arbitrariness by establishing a kurtosis threshold $\kappa_{th}$: a window with $\kappa>\kappa_{th}$ is deemed non-trivial, with scaling properties; otherwise, it reflects Gaussian activity without scaling (Fig. 2a). We allow $\kappa_{th}$ to assume any value greater than zero.
Figure 2: Improved task performance is linked to higher kurtoses. a, Distribution of kurtosis values at the final coarse-graining step (projections onto $N/16$ covariance eigenvectors) across time windows from all subjects (aggregated data). Unlike the surrogate data, the kurtosis of real data span a wider range of values. Windows with kurtosis values above (below) a tunable threshold ($\kappa_{th}$) are labelled as scaling (no-scaling) windows. b, The performance of real data is compared to data where hits and misses are shuffled (trial shuffled). c, The hit rate calculated across scaling windows is higher than both no-scaling windows and trial-shuffled data for all kurtoses exceeding the maximum value obtained for surrogate data $\kappa_{\text{max}}^{\text{surr}}$ (shaded area). Dashed line shows the kurtosis distribution. d, The relative performance gain is calculated from the difference between scaling and no-scaling curves, showing up to $24\%\pm 3\%$ increase in performance.
The experiment has three phases. During the first phase, the mice perform a task of image change recognition every few seconds and are rewarded for correct responses (hits, Fig. 2b[47]. For any value of $\kappa_{th}$ above surrogate values, the average hit rate was consistently higher when calculated over scaling windows (Methods 1.1) than over no-scaling windows (Fig. 2c). We have verified that higher kurtosis was associated to better performance (Fig. 2c), with relative performance gain reaching up to $24\%\pm 3\%$ (Fig. 2d). As a control, we confirmed that the results were better than those obtained when hits and misses were shuffled within each session (trial-shuffled data, Fig. 2b). Time windows of 10 s and 60 s were also analysed and showed similar results, associating higher kurtosis with better performance (SI Fig.4).
Figure 3: Comparison of the three phases of the experiment. a, Fraction of windows above a given kurtosis threshold ($\kappa_{th}$) in the aggregated data (complementary cumulative distribution function). The active phase remains consistently above the other phases for any choice of $\kappa_{th}$, indicating a less Gaussian-like behaviour. The shaded area corresponds to kurtoses of surrogate data. (b-e), comparison of scaling exponents ($\alpha$, $\beta$, $\mu$ and $z$) during the three phases. Bars (lines) represent the mean values (standard deviations) across sessions. Lighter shades depict results from surrogate data. Dashed lines correspond to the trivial values of exponents $\alpha$, $\beta$ and $z$. Triple asterisks (*) indicate a $p$-value $<0.001$ for the statistical comparison between two distributions utilising the Mann-Whitney U test.
In the passive phase, the same images are shown, but there is no task to be performed or reward to be gained (the lick port is removed). During this phase, the distribution of kurtoses across windows got closer to surrogate values than in the active phase. For any $\kappa_{th}$, scaling windows were more abundant in the active phase than in the passive phase (Fig. 3a). Before the passive phase, mice were presented with non-natural images, such as Gabor patches, black and white screen flashes, or a grey background [47]. In this control condition, the shift towards surrogate levels in the kurtosis distribution was even more pronounced, and scaling windows became less likely to occur (Fig. 3a). These findings indicate that the process of performing a task drives an adaptation in the dynamics of the visual cortex, directing the system toward a non-Gaussian fixed point.
We also calculated phenomenological renormalisation scaling exponents using a real-space approach to the coarse-graining method [37, 48] (Methods 1.3.2) (Fig. 3b-e). This process begins by calculating pairwise correlations from neuronal activity and progressively summing the time series of the most correlated pairs of neurons until there are none left unpaired. By iterating this process $k$ times, we obtain coarse-grained variables (or “clusters”) that are sums of $2^{k}$ spiking time series.
Near a non-Gaussian fixed point, several quantities will scale with cluster size following non-trivial power laws, namely: the mean variance (exponent $\alpha$), minus the log of the probability of silence within a cluster ($\beta$), and the mean autocorrelation characteristic time ($z$). Additionally, the covariance eigenvalues scale with their rank with exponent $\mu$ [37, 48, 43, 44, 45] (Methods 1.3.2). For the active phase, we obtained $\alpha_{1}=1.41\pm 0.05$, $\beta_{1}=0.77\pm 0.04$, $\mu_{1}=0.73\pm 0.08$ and $z_{1}=0.3\pm 0.1$ (Fig. 3b-e). These scaling exponents show minimal change when calculated for kurtoses above $\kappa_{th}$, with only a weak dependence on this threshold (SI 6). During the passive phase we found $\alpha_{2}={1.37\pm 0.04}$, $\beta_{2}=0.78\pm 0.04$, $\mu_{2}=0.61\pm 0.05$, $z_{2}=0.36\pm 0.06$, with statistically significant differences with respect to the active phase for $\alpha$, $z$ and $\mu$ (Fig. 3b-e). These values align with those reported for awake mice in a different dataset [43]. Furthermore, for non-natural stimuli the exponents were $\alpha_{3}=1.28\pm 0.03$, $\beta_{3}=0.82\pm 0.03$, $\mu_{3}=0.38\pm 0.04$ and $z_{3}=0.4\pm 0.1$, which are all different from the phases with natural images (Fig. 3b-e).
Discussion
We observed that non-Gaussian, scaling dynamics in the mouse primary visual cortex are closely linked with improved performance in a visual task, supporting the idea that critical dynamics underlie functional advantages in neural processing. These findings provide strong evidence for the critical brain hypothesis, which suggests that neural systems operate near a critical point to optimize information processing and may be essential for understanding brain function [2, 13].
The kurtosis of the coarse-grained activity distribution offers a powerful means of assessing critical dynamics, with its flexibility — particularly the tunable threshold ($\kappa_{th}$) — enabling control over which windows are classified as ‘truly’ scaling. Our findings show that high-kurtosis windows consistently correspond to improved performance, marking these periods as indicative of a near-critical state. By increasing $\kappa_{th}$ to focus on only the highest kurtosis values, performance gains in scaling windows intensify, reaching up to $24\%\pm 3\%$ over non-scaling periods (Fig. 2c-d).
Moreover, we found that the signatures of criticality in V1 are modulated by the level of task involvement. For the exact same set of images presented in the same sequence, kurtosis levels observed in the active phase drop considerably in the passive phase, simply by preventing the mouse from actively engaging with them. When presented with non-natural images, kurtoses drop even further toward trivial values. The mouse brain’s ability to self-regulate and display scaling dynamics predominantly while performing a task reinforces the idea that scaling plays a role in enhancing sensory processing.
The four scaling exponents also reflect the differences among the phases of the experiment. When animals are shown natural images and are actively performing the task, $\alpha$ and $\beta$ (“static exponents” [46]) deviate significantly from trivial values $\alpha=1$ and $\beta=1$ (Fig. 3b-c). In contrast, during a passive replay of natural images, the static exponents shift closer to trivial values, approaching this baseline even more closely when non-natural images are used as stimuli without any task involved. The “dynamic exponent” $z$ exhibited an inverse trend, diverging progressively from the trivial value $z=0$ across phases: increasing from active to passive and peaking in the non-natural image phase. This pattern is reminiscent of findings from a recently developed temporal renormalization group technique, which showed a deviation from critical behaviour during action compared to deep sleep [49]. Together, these findings provide new insights into the relationship between behaviour and criticality in the primary visual cortex, suggesting that task engagement and stimulus type modulate cortical scaling properties.
Since pioneering experiments first reported neuronal avalanches in neural systems, numerous theoretical studies have proposed that the brain’s operation near criticality confers functional advantages, enhancing processing efficiency and adaptability. Our research bridges the gap between the critical brain hypothesis and its main flagship: that criticality underlies the brain’s efficient dynamics, facilitating optimized behaviour and cognition. Furthermore, our finding that scaling properties depend on the animal’s context opens new avenues for investigating how criticality might contribute to diverse brain functions and exploring the underlying mechanisms that sustain this state across varying cognitive demands and environmental challenges.
1 Methods
1.1 Data
The data used in this study comes from the Allen Visual Behavior Neuropixels dataset [50], which includes neuronal activity recorded from the V1 region of mice using Neuropixels 1.0 probes. Each recording spans approximately 2 hours and 25 minutes. Prior to the experiment, mice were trained to lick a device upon detecting an image change in order to receive a reward. During the first hour of the experiment, they were shown a series of images under these trained conditions. Following this phase, mice were shown Gabor patches for 15 minutes, then exposed to black and white flashes, and finally shown a grey screen for 5 minutes. In the third phase, the natural images are shown a second time in the same order as before, but this time the lick apparatus is retracted. The raw electrophysiological data was pre-processed and sorted into spike trains using the Kilosort2 algorithm [47]. We selected all 62 sessions from the dataset that had more than 128 recorded units, which included both single-unit activity (SUA) and multi-unit activity (MUA). The data was then divided into these three experimental phases, binned into 50-ms intervals, and segmented into 30-s windows for analysis.
We excluded from the analysis temporal windows based on the coefficient of variance (CV)
$CV=\frac{\sigma}{\mu},$ (1)
where $\mu$ and $\sigma$ are the mean and the standard deviation of the activity across the window. Windows where $CV>5$ were excluded, as such high values are not typically observed in awake mice and likely result from the chosen window size, indicating artefacts.
1.2 Surrogate data
Throughout our analysis, we compared our results with surrogate data. This surrogate data was generated by shuffling the inter-spike intervals of each unit within each time window, effectively breaking correlations between units and producing trivial kurtosis values and exponents for comparison.
1.3 Phenomenological Renormalisation Group (PRG) analysis
1.3.1 Momentum-space renormalisation and kurtosis
The coarse graining procedure initially proposed by Bradde and Bialek [36] was applied to each of the 30-s time windows. Here we give a brief account of the method.
We start from the covariance matrix
$C_{ij}=\langle\phi_{i}\phi_{j}\rangle-\langle\phi_{i}\rangle\langle\phi_{j}\rangle,$ (2)
where $\phi_{i}$ is the time series activity of unit $i$. Taking its ranked eigenvalues ${\lambda_{1}>\lambda_{2}>\ldots>\lambda_{N}}$ and the components $u_{\mu i}$ of their respective eigenvectors, such that
$\sum^{N}_{j}C_{ij}u_{\mu j}=\lambda_{\mu}u_{\mu i}\;,$ (3)
we write the projector [36]
$\hat{P}_{ij}(N_{\mathrm{cutoff}})=\sum^{N_{\mathrm{cutoff}}}_{\mu=1}u_{\mu i}u_{\mu j}.$ (4)
Using Eq. (4) we obtain a set of variables $\psi_{i}$ for a chosen value of $N_{\mathrm{cutoff}}$
$\psi_{i}(N_{\mathrm{cutoff}})=\mathcal{Z}_{i}(N_{\mathrm{cutoff}})\sum^{N}_{j=1}\hat{P}_{ij}(N_{\mathrm{cutoff}})\left(\phi_{i}-\langle\phi_{i}\rangle\right)\;,$ (5)
where $\mathcal{Z}_{i}(N_{\mathrm{cutoff}})$ is such that $\text{var}(\psi_{i})=1$, as shown in Fig. 1.d for $N_{\mathrm{cutoff}}=N$, $N/2$, $N/4$, $N/8$, $N/16$ [48, 45].
From the distribution of $\phi_{i}$ at the final coarse graining step ($N_{\mathrm{cutoff}}=N/16$) we calculate the kurtosis $\kappa=\langle\psi^{4}\rangle/\langle\psi^{2}\rangle^{2}$ (Fig. 1.f). Non critical dynamics converge towards a Gaussian distribution ($\kappa=3$) as we reduce the number of eigenvectors involved in the construction of the projector (4), while critical dynamics does not. Note that coarse-grained distributions do not converge perfectly to a Gaussian even for surrogate data (with lowest kurtoses around $\kappa\simeq 6$), due to the finite number of bins within a window.
1.3.2 Real-space renormalisation and exponents
In order to obtain the PRG scaling exponents presented in Fig. 3.(b-e) we use a different phenomenological renormalisation process, to which we refer as real-space renormalisation [37]. In this analysis we cluster together the activities of the most correlated pair of neurons repeatedly until we exhaust all neurons:
$\sigma_{i}^{\prime(k+1)}=\sigma_{i}^{(k)}+\sigma_{j}^{(k)},$ (6)
where $j$ is the most correlated neuron to $i$. This procedure is repeated $k$ times, after which we have $N/2^{k}$ coarse-grained units (or “clusters”), each comprising the sum of $K=2^{k}$ original spiking units. We follow some quantities that are expected to scale with non-trivial power laws only close to the critical regime.
Mean variance
$M_{2}$, the variance of the activity of the clusters of size $K$, averaged over all clusters, grows with the cluster size as
$M_{2}\propto K^{\alpha}.$ (7)
Uncorrelated neurons lead to the trivial value $\alpha=1$ via the Central Limit Theorem. Perfectly correlated neurons lead to $\alpha=2$.
Silence probability
An effective free-energy can be written using the probability $P_{\mathrm{silence}}$ of finding a silent bin within each cluster:
$F=-\log(\mathcal{P_{\mathrm{silence}}})\propto K^{\beta}.$ (8)
For the trivial case of uncorrelated Poisson spiking units, one obtains $\beta=1$.
Covariance spectrum
The spectrum of eigenvalues of the covariance matrix (Eq. 2) when ordered in a descending order should follow
$\lambda=A\left(\frac{K}{\mathrm{rank}}\right)^{\mu}.$ (9)
For trivial data, this scaling is broken and $\mu$ is undefined.
Temporal correlation
Let $C(t)$ be the average autocorrelation function $C(t)$ across clusters of size $K$, from which we take the characteristic autocorrelation time $\tau_{c}$, defined by $C(\tau_{c})=1/e$. This characteristic time scales with the cluster size as
$\tau_{c}\propto K^{z}\;.$ (10)
For the trivial case of uncorrelated spike trains, one has $z=0$.
1.4 Hit rate, trial shuffling and relative performance
1.4.1 Hit rate
During the first phase of the Allen Brain Observatory Visual Behavior Neuropixels experiment, mice performed a visual recognition task. To assess the impact of criticality on performance, we subdivided the data into 30-s intervals. After applying the PRG procedure streamlined in Methods 1.3, intervals with a kurtosis above a chosen $\kappa_{th}$ were labelled as scaling windows. We then calculated the hit rate for these scaling windows by summing all correctly performed tasks (hits) within them and dividing by the total number of tasks in these intervals.
$\mathrm{Hit~{}rate(Scaling)}$ $=\frac{\mathrm{Total~{}number~{}of~{}sucessful~{}tasks(windows~{}with~{}\kappa>\kappa{th})}}{\mathrm{Total~{}number~{}of~{}tasks(windows~{}with~{}\kappa>\kappa{th})}},$ (11)
$\mathrm{Hit~{}rate(No~{}scaling)}$ $=\frac{\mathrm{Total~{}number~{}of~{}sucessful~{}tasks(windows~{}with~{}\kappa<\kappa{th})}}{\mathrm{Total~{}number~{}of~{}tasks(windows~{}with~{}\kappa<\kappa{th})}}.$ (12)
The standard error of the mean was estimated through a bootstrapping statistic method, by choosing $\mathcal{N}$ samples with replacement from the subset of time windows above (below) the kurtosis threshold, averaging and taking the standard deviation over $n=10^{5}$ repetitions.
To quantify the enhancement in performance due to scaling over no-scaling windows, we define the relative performance gain as
$\mathrm{Relative~{}performance}=\frac{\mathrm{Hit~{}rate(Scaling)}-\mathrm{Hit~{}rate(No~{}scaling)}}{\mathrm{Hit~{}rate(No~{}scaling)}}.$ (13)
1.4.2 Trial shuffled data
To control for artifacts, we generated a trial-shuffled version of the data. For each trial, we randomly selected another trial within the same session and swapped their outcomes (hit or miss, Fig. 2.b). This process preserves the total number of hits and misses within each session, as well as the total number of trials within each window.
For the trial-shuffled data, we calculated the hit rate for the scaling windows as described in Methods 1.4.1, repeating this procedure 500 times. The mean and standard deviation across these repetitions are presented in (Fig. 2.c).
1.5 Statistical procedures
All the results are expressed as mean $\pm$ standard deviation. For statistical comparisons between distributions we used Mann-Whitney U test. We use the convention for representing $p$-values: $p\geq 0.05$ (n.s.), $p<0.05$ (*), $p<0.01$ (**), $p<0.001$ (***), where “n.s.” stands for not significant. The p-value for the comparisons stated in the text are the following:
  • $\alpha_{1}$ vs $\alpha_{2}$p-value$=8\times 10^{-5}$
  • $\alpha_{2}$ vs $\alpha_{3}$p-value$=4\times 10^{-19}$,
  • $\alpha_{1}$ vs $\alpha_{3}$p-value$=8\times 10^{-21}$,
  • $\beta_{1}$ vs $\beta_{2}$p-value$=0.17$,
  • $\beta_{2}$ vs $\beta_{3}$p-value$=3\times 10^{-10}$,
  • $\beta_{1}$ vs $\beta_{3}$p-value$=3\times 10^{-13}$,
  • $z_{1}$ vs $z_{2}$p-value$=6\times 10^{-5}$,
  • $z_{2}$ vs $z_{3}$p-value$=6\times 10^{-6}$,
  • $z_{1}$ vs $z_{3}$p-value$=3\times 10^{-10}$,
  • $\mu_{1}$ vs $\mu_{2}$p-value$=4\times 10^{-13}$,
  • $\mu_{2}$ vs $\mu_{3}$p-value$=3\times 10^{-8}$,
  • $\mu_{1}$ vs $\mu_{2}$p-value$=3\times 10^{-8}$.
[…]
Supplementary information
This article contains supplementary material.
[…]
Acknowledgements The authors acknowledge funding from the Brazilian agencies CNPq (Grants No. 140660/2022-4, 308703/2022-7 and 314094/2023-7), FACEPE (Grant No. APQ-1187-1.05/22) and FINEP (Grant No. NeuroAssist 01.24.0124.00).
Declarations
Funding
The authors acknowledge funding from the Brazilian agencies CNPq (Grants No. 140660/2022-4, 308703/2022-7 and 314094/2023-7) and FACEPE (Grant No. APQ-1187-1.05/22) and FINEP (Grant No. NeuroAssist 01.24.0124.00).
Competing interests
Authors declare no competing interests.
Ethics approval and consent to participate
Not applicable.
Consent for publication
All authors consent to the publication.
Data availability
Data from all the 62 sessions used are available to download via the AllenSDK development kit at https://allensdk.readthedocs.io/.
Materials availability
Not applicable.
Code availability
The code used can be made available upon request.
Author contribution
MC and DMC designed the study. DMC provided initial code for the analysis. GGC refined the code, ran the analysis and made the plots. All authors discussed the results and wrote the paper.
Appendix A Supplementary Information
A.1 Effects of different time window lengths
Our choice of time window length to perform the analysis was made to balance two competing considerations: on the one hand, to have a sufficiently long time for the surrogate procedure to effectively shuffle the data within each window; on the other hand, to avoid an overly broad time frame that could blend different cortical dynamics.
To examine the sensitivity of our results to this choice of window length, we repeated the analyses using 10-s and 60-s windows. This allowed us to evaluate the impact of shorter and longer time windows on our performance results.
Figure 4: Relation of task performance and kurtosis for different windows sizes. (a-b), hit rates for the 10-s and 60-s windows exhibit a similar trend to those of 30-s windows, with increasing hit rates as the threshold ($\kappa_{th}$) is increased. (c-d), Relative performance follows a similar pattern, deviating from zero around $\kappa\approx 7$.
We observe that the change in window size choice does not qualitatively alter our results (Fig. 4). We consistently observe that windows with higher kurtosis are associated with improved performance.
A.2 Presence of noise units
A natural question arises regarding whether our choice to use a greedy spike sorting approach (potentially including noise units) could influence the results. To address this, we repeated the analyses using only the 43 sessions that retained more than 128 units after applying more strict sorting criteria, while maintaining the 30-s time window.
Figure 5: Relation of task performance and kurtosis for only “no noise” units. a-b We repeat the plots presented in the main text, the results of increased performance are kept.
The results remained qualitatively consistent with those observed in the initial analyses (Fig. 5). This outcome is expected, as a procedure in the spirit of the renormalisation group, which captures large-scale properties, should be resilient to the inclusion of noise units.
A.3 Exponents vs. threshold
The exponents presented in (Fig. 3) are calculated across all windows of a given session. However, it is reasonable to question this approach, as only a subset of these windows may truly exhibit scaling behaviour. To address this, we analyse how the exponents $\alpha$, $\beta$, $\mu$ and $z$ change when calculated using only the windows with $\kappa>\kappa_{th}$.
Figure 6: Exponents variation with $\kappa_{th}$. (a-d) Exponent values show only minimal variation with different $\kappa_{th}$. The light solid lines represent the surrogate data, while dashed lines show the kurtosis distribution $P(\kappa)$ of each experimental phase.
We find that the exponent values vary only slightly and remain within their respective error bars (Fig. 6). The larger fluctuations for larger kurtosis values are due to the small number of windows (note that $P(\kappa)$ has decreased considerably in those regions).
A.4 Activity kurtosis vs. coarse-grained activity kurtosis
One could wonder whether coarse-graining the neural activity is necessary to compute the kurtosis, or if examining the kurtosis of the original, non-coarse-grained activity distribution would suffice. Our analysis shows that the initial kurtosis $\kappa_{0}$ is a very poor predictor of the coarse-grained kurtosis $\kappa_{final}$ (Fig. 7).
Figure 7: Comparison of initial and final kurtoses. For each window we plot the value of the kurtosis of the activity distribution after the coarse-graining ($\kappa_{\mathrm{final}}$) against their initial kurtosis from the activity distribution ($\kappa_{0}$). There is no clear tendency in the relation between those two variables.